OpenStudy (anonymous):

George has a small coin collection. He has collected 16 gold coins, 28 silver coins, 32 copper coins, and 4 bronze coins. At a flea market, George sells 15% of his coins. Find the number of coins he sold.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Help

OpenStudy (texaschic101):

first find total number of all coins

OpenStudy (anonymous):

k um 16+28=44+32=76+4=80

OpenStudy (grace111):

Total number of coin that George has 16 + 28 + 32 + 4 = 80 coins total number of coins sold by George = 15% of 80 = 15/100 X 80 = 12 coins

OpenStudy (texaschic101):

correct...now he sold 15%...so we need to find 15% of 80 turn 15% into a decimal = 0.15 " of " means multiply (0.15)(80) = your answer...you can do the math

OpenStudy (anonymous):

k is it 12

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

tnx can u help me with more

OpenStudy (texaschic101):

not too many..I gotta go soon

OpenStudy (anonymous):

Can u help me with 2

OpenStudy (texaschic101):

I can try

OpenStudy (anonymous):

Jaime wants to know how fast she can type, so she times herself as she types a paper. The paper has 400 words in it, and Jaime types it in 8 minutes. What is her typing rate, in words per minute? A. 48 words per minute B. 50 words per minute C. 49 words per minute D. 47 words per minute

OpenStudy (texaschic101):

400/8 = 50/1 ---> 50 wds per minute

OpenStudy (anonymous):

Curtis decided to go on a road trip to Canada. On the first day of his trip, he drove for 10 hours and traveled 630 miles. At what rate did he travel on the first day, in miles per hour? A. 64 miles per hour B. 65 miles per hour C. 63 miles per hour D. 60 miles per hour

OpenStudy (texaschic101):

630/10 = 63/1 ---> 63 miles per hr

OpenStudy (anonymous):

Tnx bye!!

OpenStudy (texaschic101):

YW :)